if x+1/x= 2 find the value of 8x^3+1/x^3
Answer by lwsshak3(11628)

if x+1/x= 2 find the value of 8x^3+1/x^3
..
x+1/x= 2
x^2+1=2x
x^2-2x+1=0
(x-1)^2=0
x-1=0
x=1
..
8x^3+1/x^3
=8+1
=9